# Is Limited Participant Diversity Impeding EEG-based Machine Learning?

### v1

@@ Abstract
The application of machine learning (ML) to electroencephalography (EEG) has great potential to advance both neuroscientific research and clinical applications. However, the generalisability and robustness of EEG-based ML models often hinge on the amount and diversity of training data. It is common practice to split EEG recordings into small segments, thereby increasing the number of samples substantially compared to the number of individual recordings or participants. We conceptualise this as a multi-level data generation process and investigate the scaling behaviour of model performance with respect to the overall sample size and the participant diversity through large-scale empirical studies. We then use the same framework to investigate the effectiveness of different ML strategies designed to address limited data problems: data augmentations and self-supervised learning. Our findings show that model performance scaling can be severely constrained by participant distribution shifts and provide actionable guidance for data collection and ML research.
1 Main
@@
A century after the first human electroencephalogram was recorded in 1924, Electroencephalography (EEG) is in many ways still the best available technique for the direct and non-invasive measurement of brain activity. In particular, with newer generations of dry-electrode recording hardware, it has a unique set of desirable properties, including its low cost and patient burden, portability, and high temporal resolution. Given the complexity and high dimensionality of EEG signals, machine learning (ML) has shown considerable promise in advancing neuroscientific research and clinical applications [gemein_machine-learning-based_2020, wu_electroencephalographic_2020, jing_development_2023, mitchell_crow_could_2024]. However, the successful deployment of ML in EEG-based applications is not without challenges. A fundamental obstacle is the limited availability of high-quality EEG data, which is often sparse, noisy, and costly to collect.
@@
While prior works have acknowledged limited data as a challenge for EEG-based machine learning [roy_deep_2019, rommel_data_2022], systematic studies are currently lacking. Furthermore, an aspect that has largely been ignored is the hierarchical nature of the data generation process relevant to EEG machine learning work. EEG-based ML models rarely operate on full recordings; samples usually correspond to more manageable segments on the order of a few seconds, extracted from recordings through sliding-window approaches [roy_deep_2019]. As a result, the overall sample size is usually considerably larger than the number of individual participants in the dataset. Many applications, particularly in a medical context, require models to generalise to new participants – e.g., a diagnostic model is of little use if it can only diagnose patients in the training data. Consequently, if there is a significant distribution shift between participants, low participant diversity [^1] In this study, ‘participant diversity’ refers to the number of individuals contributing data, with higher diversity corresponding to a greater number of unique participants. could impede model performance even with a large overall sample size. Understanding the relevance of participant diversity is critical to guide data collection and future EEG-based machine learning work. If models generalise well across participants, data could be collected from smaller cohorts, significantly reducing costs and logistical challenges. Conversely, if participant diversity is crucial, this motivates not only higher participant counts during data collection, but also the development of machine learning strategies that explicitly address participant distribution shifts.
@@
To address these challenges, we first formalise a hierarchical data generation process that captures the structure of EEG data used to train machine learning models. Furthermore, we systematically investigate the data scaling behaviour of different EEG machine learning models (TCN, mAtt, LaBraM) across a range of large datasets (TUAB, CAUEEG, PhysioNet), each with more than 1,000 participants, and three tasks (EEG normality prediction, dementia diagnosis, sleep staging). Crucially, we control both the number of participants in the training data and the amount of data per participant, allowing us to disentangle the impact of participant diversity from the overall sample size.
Beyond the collection of larger datasets, several works of research have proposed strategies to deal with limited data by making better use of available data or by transferring representations learned on other (sometimes unlabelled) datasets. Inspired by their success in computer vision, one line of research has focused on the development of data augmentations [rommel_data_2022, rommel_cadda_2022, george_data_2022, mohsenvand_contrastive_2020], which artificially inflate the dataset size through the application of label-preserving transformations to available samples [shorten_survey_2019]. Often these transformations are designed based on prior knowledge about desirable invariances, e.g. prominent examples in computer vision like cropping or translation encourage invariance to object positions within the image [shorten_survey_2019, ericsson_self-supervised_2022]. Moreover, the success of large language models has sparked a lot of interest in pre-training strategies based on masking and the reconstruction of raw data or latent features [chien_maeeg_2022, cui_neuro-gpt_2023, jiang_large_2023, song_eeg_2022]. Several of these works have also adopted the notion of foundation models [bommasani_opportunities_2022]. These studies have laid valuable groundwork, often demonstrating effectiveness on smaller datasets or for particular prediction tasks. However, understanding how these approaches perform across a broader range of data settings can provide deeper insights into their generalisability and scalability.
We assess the utility of existing data augmentations (AmplitudeScaling, FrequencyShift, PhaseRandomisation) and a pre-trained foundation model (LaBraM) for varying amounts of (downstream) training data. Specifically, we quantify their effectiveness in data-limited regimes, whether they continue to provide value as the dataset size increases, and how they address specific limitations, such as low participant diversity versus limited data per participant.
In summary, we offer the following contributions:
@@
  • We formalise the data generation process for EEG-based machine learning and use recent, relevant advances in statistical learning theory to inform our experiment designs.
  • We empirically study the scaling behaviour and the relevance of participant distribution shifts across different models, tasks, and datasets.
  • We assess the utility of existing data augmentations and a pre-training strategy in improving model performance in different data regimes.
@@ 2 Results
2.1 Data Generation Process and Experiment Overview
We model datasets used for EEG-based machine learning by the hierarchical data generation process illustrated in Figure 1A. A dataset comprises recordings from $n$ different participants, which are in turn segmented into $m$ distinct samples per participant. As a result, the overall sample size is usually substantially larger than the number of individual participants from which the samples are obtained. Formally, we can denote the sample features by $X_{ij}$ where $i=1,...,n$ indexes the participants, and $j=1,...,m$ indexes the samples for participant $i$, and with corresponding ground-truth labels denoted by $Y_{ij}$. For each participant, these samples follow a participant-specific distribution $P_{i}$, which is itself drawn from a population-level distribution $Q$.
Consider a loss function, $\ell(\hat{Y},Y)$, measuring the quality of a prediction, $\hat{Y}=f(X)$, provided by some model, $f$. This could be, e.g., the zero-one loss or the cross entropy loss, and need not be the loss actually used to train a model. We can define the empirical risk (measured on the training set) and population risk as
$r(f):=\frac{1}{mn}\sum_{i=1}^{n}\sum_{j=1}^{m}\ell(f(X_{ij}),Y_{ij})$ (1)
and
$R(f)=\mathbb{E}_{X,Y}[\ell(f(X),Y)],$ (2)
respectively. In the classic learning setting, where each $P_{i}$ is assumed to be the same, it is well established in the statistical learning theory literature that the gap between empirical risk and population risk (i.e., the extent to which a model will overfit) scales as $\tilde{\Theta}(1/(mn)^{p})$, where $p\in\{\frac{1}{2},1\}$ [hanneke_revisiting_2024]. The value of $p$ will be $\frac{1}{2}$ unless some assumptions are satisfied relating to the suitability of the model architecture and the amount of information the features contain about the labels [hanneke_optimal_2016]. The two-level data generation process we consider is comparatively less studied, but it is known that with minimal assumptions about model choices and feature information content, the gap will scale as $\tilde{\Theta}(1/n^{\frac{1}{2}}+1/(mn)^{\frac{1}{2}})$ [gouk_limitations_2024]. We include the second term, which is asymptotically dominated by the first, because if the distribution shift is small, or one is able to select a model architecture with substantial robustness to distribution shifts, then the second term can be larger in the non-asymptotic regime [gouk_limitations_2024]. This analysis tells us that in the presence of a substantial distribution shift, the extent to which one should expect to overfit is governed primarily by the number of participants, rather than the amount of data available per participant, unless one is able to design a model architecture with the appropriate invariances.
Figure 1: (A) Multi-level EEG data distribution. A dataset contains EEG recordings from multiple participants, which are usually further divided into smaller segments to train machine learning models. (B) Grid visualising the subsampling of training datasets controlling for both the number of participants and the number of segments per participant. (C) The datasets, tasks, and models used in the experiments.
We conducted data scaling experiments to empirically assess the dependence of ML model performance on the participant count $n$ and the number of samples per participant $m$ across three large datasets and prediction tasks (normality prediction on TUAB – binary classification, dementia diagnosis on CAUEEG – three-way classification, and sleep staging on PhysioNet – five-way classification; see Section 4.2). Starting from the full datasets, we randomly subsampled training datasets with a fixed number of participants and overall sample size as visualised by the grid depicted in Figure 1B. Different ML models (TCN – a convolutional neural network, mAtt – a geometric deep learning approach, and LaBraM – a transformer architecture; see Section 4.3) were then trained on the subsampled training data. Early stopping was used to prevent overfitting with limited amounts of training data, while still allowing training to converge for larger training datasets. See Section 4.1 for more details on the experimental setup and model training.
2.2 Scaling Behaviour of Model Performance
Figure 2 depicts model accuracies for a fixed number of segments per participant ($m=40$) and shows how performance scaled with increasing sample size (specifically participant count) across the different datasets. While the shapes of the curves align with expectations from a machine learning perspective, it is informative to look at the magnitude of improvements and the participant counts at which they started to diminish.

### v2

@@ Abstract
The application of machine learning (ML) to electroencephalography (EEG) has great potential to advance both neuroscientific research and clinical applications. However, the generalisability and robustness of EEG-based ML models often hinge on the amount and diversity of training data. It is common practice to split EEG recordings into small segments, thereby increasing the number of samples substantially compared to the number of individual recordings or participants. We conceptualise this as a multi-level data generation process and investigate the scaling behaviour of model performance with respect to the overall sample size and the participant diversity through large-scale empirical studies. We then use the same framework to investigate the effectiveness of different ML strategies designed to address limited data problems: data augmentations and self-supervised learning. Our findings show that model performance scaling can be severely constrained by participant distribution shifts and provide actionable guidance for data collection and ML research. The code for our experiments is publicly available online. [^1] https://github.com/bomatter/participant-diversity-paper
1 Introduction
@@
A century after the first human electroencephalogram was recorded in 1924, Electroencephalography (EEG) is in many ways still the best available technique for the direct and non-invasive measurement of brain activity. In particular, with newer generations of dry-electrode recording hardware, it has a unique set of desirable properties, including its low cost and patient burden, portability, and high temporal resolution. Given the complexity and high dimensionality of EEG signals, machine learning (ML) has shown considerable promise in advancing neuroscientific research and clinical applications [1, 2, 3, 4]. However, the successful deployment of ML in EEG-based applications is not without challenges. A fundamental obstacle is the limited availability of high-quality EEG data, which is often sparse, noisy, and costly to collect.
@@
While prior works have acknowledged limited data as a challenge for EEG-based machine learning [5, 6], systematic studies are currently lacking. In particular, an aspect that has largely been ignored is the hierarchical nature of the data generation process relevant to EEG machine learning work. EEG-based ML models rarely operate on full recordings; samples usually correspond to more manageable segments on the order of a few seconds, extracted from recordings through sliding-window approaches [5]. As a result, the overall sample size is usually considerably larger than the number of individual participants in the dataset. Many applications, particularly in a medical context, require models to generalise to new participants—e.g., a diagnostic model is of little use if it can only diagnose patients in the training data. Consequently, if there is a significant distribution shift between participants, low participant diversity [^2] In this study, ‘participant diversity’ refers to the number of individuals contributing data to the training dataset, with higher diversity corresponding to a greater number of unique participants. could impede model performance even with a large overall sample size. Understanding the relevance of participant diversity is critical to guide data collection and future machine learning work. If models generalise well across participants, data could be collected from smaller cohorts, significantly reducing costs and logistical challenges. Conversely, if participant diversity is crucial, this motivates not only higher participant counts during data collection, but also the development of machine learning strategies that explicitly address participant distribution shifts.
@@
We address these challenges by systematically investigating the data scaling behaviour of different EEG machine learning models (TCN, mAtt, LaBraM) across a range of large datasets (TUAB, CAUEEG, PhysioNet), each with more than 1,000 participants, and three tasks (EEG normality prediction, dementia diagnosis, sleep staging). Crucially, we control both the number of participants in the training data and the amount of data per participant, allowing us to disentangle the impact of participant diversity from the overall sample size. Furthermore, we assess the utility of existing data augmentations (AmplitudeScaling, FrequencyShift, PhaseRandomisation) and a pre-trained foundation model (LaBraM) for varying amounts of (downstream) training data. Specifically, we quantify their effectiveness in data-limited regimes, whether they continue to provide value as the dataset size increases, and how they address specific limitations, such as low participant diversity versus limited data per participant.
@@
In summary, we formalise the data generation process for EEG-based machine learning and use recent, relevant advances in statistical learning theory to inform our experiment designs and answer the following research questions:
  • Q1: Does a lack of participants in EEG datasets substantially limit the performance of machine learning models?
  • Q2: Is collecting more data per participant a reliable way to compensate for a lack of participants?
  • Q3: Can contemporary EEG data augmentations effectively increase performance across different data regimes, including limited participant counts or data per participant?
  • Q4: Can self-supervised pre-training effectively increase performance across different data regimes, including limited participant counts or data per participant?
@@
2 Related Work
EEG Data Scaling Studies[7] have studied scaling laws for image decoding from brain activity measured with EEG, MEG, or fMRI [7]. A critical difference to this work is that they trained participant-specific models, where data from the same participant is used during training and test time and models do not have to generalise to unseen participants. Our work is focused on settings where generalisation to new participants is crucial, such as applications like diagnosis or prognosis. [8] specifically investigated the scaling behaviour of convolutional neural networks for EEG normality prediction on the TUAB dataset and presented scaling laws for model and data size [9]. While their experiments controlled for data size in terms of the number of recordings, we aim to disentangle the overall sample size, controlling for both the number of participants and the amount of data per participant. This allows us to study the relevance of participant distribution shifts and the need for sufficient participant diversity in datasets used for model training. Our conclusions are also more broadly applicable, as we study different tasks, datasets, and more diverse model architectures.
EEG data augmentations  Several works have proposed and employed EEG data augmentations [6, 10, 11, 12, 13, 14]. [*]mohsenvand_contrastive_2020 introduced several augmentations in the context of a contrastive learning approach for EEG [12]. [6] benchmarked various EEG data augmentations on a sleep staging dataset with 78 participants and a motor imagery task with 9 participants [10]. We build on their work, including two promising augmentations from [6] and one from [*]mohsenvand_contrastive_2020 in our experiments. Our work is different in that we specifically investigate the utility of data augmentations in dealing with limited participant counts and limited data per participant. Furthermore, our experiments were conducted on much larger datasets with over 1,000 participants each, allowing us to study the impact of augmentations across a broader range of data regimes. Finally, their experiments were limited to a single model per task, whereas we benchmarked all augmentations with different model architectures for each task and observed differential effects.
Self-supervised pre-training  Recently, a lot of work has also been published on pre-trained EEG (foundation) models [15, 16, 17, 18, 19], most of which is based on transformer models operating on patches that correspond to channel-wise EEG segments and masked token prediction. Our experiments include LaBraM as a state-of-the-art example of these approaches and provides an independent evaluation of the utility of self-supervised pre-training across different data regimes. Furthermore, our aim was again to disentangle the effect of participant counts vs overall sample size.
3 Methods
We model datasets used for EEG-based machine learning by the hierarchical data generation process illustrated in Figure˜1A. A dataset comprises recordings from $n$ different participants, which are in turn segmented into $m$ distinct samples per participant. As a result, the overall sample size is usually substantially larger than the number of individual participants from which the samples are obtained. Formally, we can denote the sample features by $X_{ij}$ where i=1,,ni=1,...,n indexes the participants, and j=1,,mj=1,...,m indexes the samples for participant $i$, and with corresponding ground-truth labels denoted by $Y_{ij}$. For each participant, these samples follow a participant-specific distribution $P_{i}$, which is itself drawn from a population-level distribution $Q$. We use a loss function, $\ell(\hat{Y},Y)$, to measure the quality of a prediction, $\hat{Y}=f(X)$, provided by some model, $f$. This could be, e.g., the zero-one loss or the cross entropy loss, and need not be the loss actually used to train a model. We can define the empirical risk (measured on the training set) and population risk, respectively, as
$r(f):=\frac{1}{mn}\sum_{i=1}^{n}\sum_{j=1}^{m}\ell(f(X_{ij}),Y_{ij})\quad\text{and}\quad R(f)=\mathbb{E}_{X,Y}[\ell(f(X),Y)].$ (1)
In the classic learning setting, where each $P_{i}$ is assumed to be the same, it is well established in the statistical learning theory literature that the gap between empirical risk and population risk (i.e., the extent to which a model will overfit) scales as $\tilde{\Theta}(1/(mn)^{p})$, where $p\in\{\frac{1}{2},1\}$ [20]. The value of $p$ will be $\frac{1}{2}$ unless some assumptions are satisfied relating to the suitability of the model architecture and the amount of information the features contain about the labels [21]. The two-level data generation process we consider is comparatively less studied, but it is known that with minimal assumptions about model choices and feature information content, the gap will scale as $\tilde{\Theta}(1/n^{\frac{1}{2}}+1/(mn)^{\frac{1}{2}})$ [22]. We include the second term, which is asymptotically dominated by the first, because if the distribution shift is small, or one is able to select a model architecture with substantial robustness to distribution shifts, then the second term can be larger in the non-asymptotic regime [22]. This analysis tells us that in the presence of a substantial distribution shift, the extent to which one should expect to overfit is governed primarily by the number of participants, rather than the amount of data available per participant, unless one is able to design a model architecture with the appropriate invariances. Under this framework, we address the four research questions outlined in the introduction. We assess the relevance and magnitude of participant distribution shifts through the scaling behaviour of model performance (Q1 & Q2) and the effectiveness of augmentations (Q3) and self-supervised pre-training (Q4) to improve performance by endowing models with useful invariances or boosting effective participant counts.
Figure 1: (A) Multi-level EEG data distribution. A dataset contains EEG recordings from multiple participants, which are divided into smaller segments to train machine learning models. (B) Grid visualising the subsampling of training datasets controlling for both the number of participants and the number of segments per participant. (C) The datasets, tasks, and models used in the experiments.
3.1 Experimental Design and Model Training
We conducted data scaling experiments to empirically assess the dependence of ML model performance on the participant count $n$ and the number of samples per participant $m$ across three large datasets and prediction tasks (see Section˜3.2). Each dataset was split into train, validation, and test splits with no participant overlap. Subsequently, we randomly subsampled training datasets with a fixed number of participants and overall sample size from the train split as visualised by the grid depicted in Figure˜1B. Different ML models (see Section˜3.4) were then trained on the subsampled training data (see Section˜A.1 for details and hyperparameters). Early stopping (on the fixed, shared validation set) was used to prevent overfitting with limited amounts of training data, while still allowing training to converge for larger training datasets. To obtain uncertainty estimates, each trial—i.e. each combination of participant count and number of samples per participant—was repeated multiple times with different random seeds for the training data subsampling. Finally, all trained models were evaluated on the same test data. To further study different strategies to deal with limited data, we repeated the same procedure for each studied method designed to address a lack of data (data augmentation: see Section˜3.5, self-supervised learning: see Section˜3.6).
3.2 Data and Tasks
TUAB  The Temple University Hospital Abnormal (TUAB) EEG Corpus [23, 24], a subset of the TUH EEG Corpus [25], was collected in a hospital environment and is provided with labels about the normality of the EEG according to standardised criteria [26]. The full dataset contains more than 2300 participants with most recording durations in the range of 15-25 minutes. For the current study, we used version 3.0.1 of the dataset. Data access can be requested from the dataset authors.
CAUEEG  The Chung-Ang University Hospital EEG (CAUEEG) dataset [27] comprises data from 1379 participants with most recording durations in the range of 5 - 20 minutes. In the current study, we used a subset of data from 1122 participants for which diagnostic labels for the categories normal, mild cognitive impairment (MCI), or dementia were available. Diagnostic labels were assigned by neurologists based on a neuropsychological examination. Data access for academic and research purposes can be requested from the dataset authors.
PhysioNet  The dataset referred to as PhysioNet in this study corresponds to the data used in the PhysioNet/Computing in Cardiology Challenge 2018 [28, 29]. While the official challenge was focused on the classification of arousal regions, we focused on sleep staging here. The dataset comprises polysomnographic (PSG) recordings along with sleep stage annotations for non-overlapping 30-second epochs following AASM standards. While the PSG recordings contained additional measurements such as electrooculography (EOG), electromyography (EMG), or airflow and oxygen saturation, only the six EEG channels were used in this study. Furthermore, we only used the training data of version 1.0.0 of the dataset, as annotations for the official test split were not made publicly available. The data is available under the Open Data Commons Attribution License (ODC-By) v1.0.
3.3 Preprocessing
Preprocessing of the data comprised the following steps: Band-pass filtering to 1-50 Hz and resampling to a sampling frequency of 100 Hz, cropping (to 15 minutes for TUAB, omitting the first 30 seconds due to frequent strong artefacts at the beginning of recordings; to 5 minutes for CAUEEG), selection of a fixed channel subset (the 19 10-20 channels for TUAB and CAUEEG; the 6 EEG channels for PhysioNet), re-referencing to average reference, and segmentation into non-overlapping epochs (2 seconds for TUAB and CAUEEG; 30 seconds for PhysioNet).
For compatibility with the pre-trained LaBraM model, the following adjustments were made for the training and fine-tuning of LaBraM: amplitudes scaled to units of 0.1 mV, sampling frequency of 200 Hz, filtering to 0.1-75 Hz with a notch filter at 50 Hz (see [15]). We note that the notch filter was not matched to the line noise of the data in the original publication (e.g. LaBraM was also evaluated on TUAB, where line noise is at 60 Hz). For compatibility with the pre-trained weights, we did not correct this potentially suboptimal setting. Furthermore, since LaBraM only supports segment lengths up to 16 seconds without modification (due to the implementation of the temporal positional encodings), we used 15 second epochs instead of 30 second epochs for PhysioNet.
3.4 Models
We selected three influential and widely used models with distinct architectures to obtain a diverse perspective on EEG model behaviour and reduce the risk of model-specific effects.
TCN  The temporal convolutional network (TCN) architecture was originally proposed by [30] [30] and has since been adapted and widely used for EEG data [1]. The model consists of a stack of 1D causal convolution layers. We used the braindecode implementation [31] with the following hyperparameters: n_blocks=4, n_filters=64, kernel_size=5, drop_prob=0.
mAtt  The manifold attention network (mAtt) is a geometric deep learning model developed for EEG data [32]. Briefly, the architecture starts with two convolutional layers and then encodes segments in the input EEG into a sequence of covariance matrices. The diagonals in these covariance matrices correspond to the signal power in each channel. The architecture then makes use of the fact that the covariance matrices are symmetric positive definite (SPD) through a manifold attention module, where instead of dot product attention, an approximation of the Riemannian distance is used. The output is then transformed back to Euclidian space and passed through a fully connected layer to get classifications. We adapted the code from the official mAtt repository (https://github.com/CECNL/MAtt) and used 100 kernels in conv1, 50 kernels in conv2 with a kernel size of 11, and an embedding size of 25 for the SPD transforms.
LaBraM  The Large Brain Model (LaBraM) architecture was designed by [15] [15]. The model operates on patches corresponding to channel-wise time segments. Patches are first passed through a small convolutional encoder and spatial and temporal positional encodings are added. The resulting tokens are then processed through a stack of transformer encoder layers and average pooling is used to aggregate the output tokens, followed by a classification head. We used the LaBraM Base model with the recommended default hyperparameters.
3.5 Augmentations
We selected three promising augmentations based on prior work (see Section˜A.2 for details).
Amplitude Scaling  The AmplitudeScaling data augmentation scales (multiplies) each EEG channel by a random factor $a\in[0.5,\ 2]$. This augmentation was proposed by [12] and used for contrastive learning [12]. They consulted four neurologists to identify augmentations that would not change the interpretation of the data along with recommended ranges for the parameters (min and max scaling factor in this case; we used the same settings in our experiments).
Frequency Shift  FrequencyShift was introduced in [10] and further evaluated in [6]. This augmentation introduces a small shift in the frequency domain, i.e. the power spectrum is shifted by some random frequency $\Delta f$. The same shift is applied to all EEG channels. [6] tuned the hyperparameter defining the maximum shift for a sleep staging task and achieved the best results with $\Delta f_{max}=0.3\,\text{Hz}$. We used the same parameter in our experiments.
Phase Randomisation[13] introduced the randomisation of Fourier transform phases for data augmentation under the name of "FT surrogates" [13]. We refer to this augmentation as PhaseRandomisation. [6] evaluated different hyperparameters for the maximum phase shift and observed the best results with high values [6]. We adopt the maximum range ($\Delta\varphi_{max}=2\pi$), which corresponds to complete randomisation of the phase information. We also sample different phases for each EEG channel, the setting that [6] used for their sleep staging experiments.
3.6 Pre-training
The LaBraM model (see Section˜3.4) was developed as a foundation model for EEG data and pre-trained to predict discrete tokens, previously learned by a separate tokeniser that was itself trained to predict the Fourier spectrum (see the original publication for details [15]). The model was pre-trained in an unsupervised manner on a collection of 16 datasets with a total of around 500 participants, and the weights for the smallest version (LaBraM Base) were made publicly available by the authors. We used these weights for experiments with the pre-trained LaBraM model.
The model was then fine-tuned using the labelled training data of the corresponding task. We fine-tuned the entire architecture without freezing any layers and determined the learning rate through hyperparameter optimisation. Specifically, we evaluated a grid of learning rates [1e-7, 5e-7, 1e-6, …, 1e-2] for each dataset (TUAB, CAUEEG, PhysioNet) using the respective training and validation splits. To ensure that performance differences between the pre-trained and baseline (trained from scratch) models can be attributed to pre-training rather than tuning of the learning rate, we also tuned the learning rate for the baseline in the same way. Table˜1 shows the selected learning rates.
@@ 4 Results
4.1 Scaling the Number of Participants
Q1: Does a lack of participants in EEG datasets substantially limit the performance of machine learning models?